OpenStudy (anonymous):

7 (f-1) = 45 Step by step:)

OpenStudy (amorfide):

first of all you must multiply out of the brackets so you will multiply everything inside the brackets by 7 7(f-1)=45 (7xf) + (7x-1)=45 7f-7=45 now you must get the unknown on one side and the constants on the other well we have a constant on the right already so we will get the negative 7 onto the right hand side to do this you will do the opposite of the sign infront of the number 7, since this is a minus sign, you will add 7 onto the left side, but if you do something to the left side you must also do it to the right side so you now have 7f-7+7=45+7 this simplifies to, 7f=52 now you have the value for 7f you want f on its own, so you want 1f you have 7 times f, so you must do the opposite, which is divide you will divide by 7 on the left hand side, but again you must do the same to the right hand side you now have 7f/7 = 52/7 7/7=1 so you are left with f you now have f=52/7

OpenStudy (anonymous):

Thanks:) I appreciate it:)